# Feature Fusion from Head to Tail: an Extreme Augmenting Strategy for Long-Tailed Visual Recognition

### v1

@@ Abstract
The imbalanced distribution of long-tailed data poses a challenge for deep neural networks, as models tend to prioritize correctly classifying head classes over others so that perform poorly on tail classes. The lack of semantics for tail classes is one of the key factors contributing to their low recognition accuracy. To rectify this issue, we propose to augment tail classes by borrowing the diverse semantic information from head classes, referred to as head-to-tail fusion (H2T). We randomly replace a portion of the feature maps of the tail class with those of the head class. The fused feature map can effectively enhance the diversity of tail classes by incorporating features from head classes that are relevant to them. The proposed method is easy to implement due to its additive fusion module, making it highly compatible with existing long-tail recognition methods for further performance boosting. Extensive experiments on various long-tailed benchmarks demonstrate the effectiveness of the proposed H2T. The source code is temporarily available at https://github.com/Keke921/H2T.
@@ 1 Introduction
(a) Comparison of class boundaries. The orange line represents the decision boundary. $\bullet$ and $\blacktriangle$ denotes head and tail class samples, respectively.
(b) Frequency of predictive labels for tail classes samples. A large number of tail class samples are incorrectly recognized as head classes.
Figure 1: The relationship between tail and head classes.
Deep neural networks (DNNs) have shown remarkable capabilities in diverse visual recognition tasks [16, 23, 6, 60, 34], yet their performance heavily relies on data that evenly distribute across categories. In contrast, real-world data typically exhibit long-tailed distributions [37, 33, 54], which has become one of the bottlenecks limiting the advancement of DNNs and hinders their broader practical applications. For example, misdiagnosis of rare diseases by computer-aided diagnosis [45, 2] may result in irreversible consequences in severe cases. Recently, a large number of methods have been proposed to address the class imbalance problem in long-tailed data. Class-balancing methods, including re-sampling [5, 18, 39, 20, 35] and re-weighting [17, 12, 4, 30, 15], are the most straightforward approaches to tackle class imbalance. Previous class-balancing methods, however, improve the overall performance but are usually at the risk of overfitting to the minority classes or the loss of information [52, 36, 20]. To overcome these limitations, recent studies have proposed alternative techniques such as re-margining methods [4, 27, 26] and ensembling learning [47, 24, 57, 10]. These methods can generate more robust predictions, but they do not consider enhancing the sample diversity of tail classes. However, insufficient samples and a lack of tail class semantics will also lead to a low recognition rate for tail classes. Data augmentation methods [46, 36, 20, 7, 28] are thus proposed to effectively enrich tail classes. While these methods have shown state-of-the-art performance, they generally aim to obtain a more representative and separable feature space and do not dig into the potential of the classifier in the obtained feature space.
This paper thereby aims to unleash the power of the obtained feature space to get a better classifier. Take the binary case in the obtained feature space as an example, the semantic space can be divided into three areas: head-like, tail-like, and hard negative. The former two areas represent the main distribution spaces of the head and tail classes, respectively, while the hard negative area includes samples that are difficult to distinguish between the two classes. In long-tail data, the head class is usually sufficiently sampled, and its semantic area is fully occupied. Besides, some samples of head class are distributed in the hard negative area. However, only a few samples are often distributed in the tail class semantic regions due to insufficient tail class sampling, resulting in a decision boundary that is limited to the semantic area of the tail class, as shown in the upper of Figure 1a.
@@
To address the issue of the lack of tail class semantics, we propose a direct and effective solution, named head-to-tail fusion (H2T), which borrows partial of the semantics from the head class and applies them to the tail class. Our approach is based on the assumption that predictions on rare instances are easily affected by objects that appear frequently. Figure 1b shows a simple experiment [^1] The experiment was conducted on CIFAR10-LT with an imbalance ratio of 100. The backbone network is ResNet-32 which was trained using cross-entropy loss. Classes 7-9 (0-indexed) are tail classes.for this assumption. The label frequency of predicting tail classes into each class is presented. We can observe that a large number of tail classes are misclassified as head classes by the model, which may indicate that the hard negative area for tail classes is filled with head class samples. Therefore, there is relatively more common closet semantics between head and tail classes. By transferring such semantics, we can effectively fill the tail semantic area and improve the recognition rate of the tail class samples, as shown in the lower of Figure 1a. On the other hand, by obtaining partial specific semantics (not common semantics) of the head class, the tail class samples can occupy the hard negative semantic region, which forces the classification boundary to move from the tail semantic region to the hard negative semantic region. This approach brings the classification boundary closer to the theoretical decision boundary, and improves the classification performance for tail classes. To streamline the fusion operation, we design a simple and easy-to-implement strategy that takes full use of the obtained features without retraining the backbone. Specifically, we randomly substitute several channels of the feature maps of the class-balanced data with that of the head-biased data. This makes the feature maps of tail classes with a high probability of being fused with head classes. H2T can be flexibly integrated with other methods, as it does not alter the structure of the backbone as well as any other technical details. The main contributions of this paper are summarized as follows:
@@
  • We propose a novel H2T that borrows information from head classes and applies it to augment tail classes without requiring additional data or network parameters, which can enrich tail classes and release the potential of obtained feature space.
@@
  • We devise a simple fusion strategy that partially substitutes the feature maps of an instance-wise sampled data branch with those of a class-balanced sampled branch. The adoption of this approach enables H2T to be implemented with ease and without requiring any modifications to the original model structure.
@@
  • Considerable performance improvement is achieved in both single and multi-expert models by integrating the H2T module into the original structure. Extensive experiments on popular benchmarks validate the efficacy of the proposed method.
@@ 2 Related Work
This section mainly makes an overview of methods based on data augmentation that is the most relevant regime.
@@
Classical Augmentation: Classical augmentation methods including flip, rotate, crop, padding, and color jittering etc. [41, 13] have been widely applied in the process of deep model training. On long-tailed data, these methods can improve model robustness and prevent overfitting to specific input patterns to a certain extent. Recently, AutoAugment [8], Fast AutoAugment [29], Population Based Augmentation [14] and Randaugment [9], etc. have been proposed to automatically determine the best data augmentation strategy for a dataset. These methods have been shown to increase classification accuracy for DNNs and demonstrated their efficacy on long-tailed data [38, 11, 52, 25].
@@
MixUp Based Augmentation: MixUp [53] is an effective data augmentation method that linearly combines a pair of row images and their labels. This method adds complexity control to the uncovered space in the data space through linearly interpolating discrete sample points, which reduces the model generalization error. Numerous studies [55, 56, 27] have experimentally proven that MixUp can significantly improve long-tail recognition performance. CutMix [50] is a variant of MixUp that augments data by cutting out a patch from the input and pasting it onto another image in the training set. The ground truth labels are combined in proportion to the patch areas. Park et al. [36] proposes to mix the foreground patches from tail classes with the background images from head classes via CutMix, enabling the transfer of context-rich background information from head to tail.
@@
New Images Based Augmentation: In addition to the aforementioned methods, researchers have proposed alternative approaches that generate new informative samples for tail classes. One such method is M2m [20], which translates samples from head classes to the tail using a pre-trained auxiliary classifier. Wang et al. [46] utilize the learned encoded variation information from head classes to synthesize images for tail classes. Zada et al. [52] directly exploit pure noise images as tail class samples. Zhang et al. [55] utilize class activation maps (CAM) [58] to separate image foreground and background, followed by augmenting the foreground through flipping, rotating and scaling, etc. The new samples are then generated by overlaying the augmented foreground on the unchanged background.
@@
Feature Level Augmentation: Data augmentation can also be performed in feature space. For example, Manifold MixUp [44] conducts the linear interpolation on the features of input images, which has been shown to yield better results in long-tail learning [55, 57]. Chu et al. [7] use CAM to decompose features into class-generic and class-specific components. Tail classes are then augmented by combining their class-specific components with the class-generic components. Another promising technique is to transfer knowledge from the head classes to the tail classes. For example, Yin et al. [49] and Liu et al. [32] enrich tail classes with the intra-class variance of head classes to balance the class distributions in feature space. Recently, meta-learning-based augmentation has been leveraged to address class imbalance. Liu et al. [31] design a meta-embedding that uses a memory bank to enrich tail classes. MetaSAug [28] learns class-wise covariance matrices by minimizing the loss on a small balanced validation set and subsequently uses the learned matrices to semantically augment tail classes.
@@ 3.1 Preliminaries
We formally define the basic notations used in this paper before going into detail about our proposed method.
For dataset: We use $\{x,y\}$ to represent one input image and its corresponding label. The total number of classes is denoted by $C$, and the training set includes $N$ samples. Thus, we have $y\in\{0,1,\cdots,C-1\}$. Suppose class $i$ has $n_{i}$ training samples. Then $N=\sum_{i}n_{i}$. For simplicity, we suppose $n_{0}\geq n_{1}\geq\cdots\geq n_{C-1}$
For backbone network: Feed $x$ into the DNN backbone, we can obtain its feature maps before the last pooling layer, denoted as $\mathcal{F}=\left[F_{1},F_{2},\cdots,F_{d}\right]\in\mathbb{R}^{w_{F}\times h_{F}\times d}$, where $w_{F}$ and $h_{F}$ represent the width and height of the feature map, respectively. $d$ is the dimension of features in embedding space. The representation after the last pooling layer is ${f}\in\mathbb{R}^{d}$. The weight of the linear classifier is represented as $\textbf{W}=[w_{1},w_{2},\cdots,w_{c}]\in\mathbb{R}^{d\times C}$, where $w_{i}$ represents the classifier weight for class $i$. We use the subscripts $h,m$ and $t$ to indicate head, medium and tail classes, respectively. $z_{i}=w_{i}^{T}f$ represents the predicted logit of class $i$, where the subscript $i=y$ denotes the target logit and $i\neq y$ denotes the non-target logit.
@@ 3.2 Motivation
We begin by analyzing the compression of the tail distribution caused by imbalances in class sizes. The standard training method for DNNs typically involves either empirical risk minimization (ERM) [43, 51] or structural risk minimization (SRM) [22]. Since the extra item in SRM is a function of model parameters, we can analyze the effect of data on DNNs without loss of generality by focusing solely on ERM. To facilitate this analysis, we partition the data by class and use the superscript $i$ to indicate their corresponding labels. Formally, ERM can be expressed as:
$\mathcal{R}_{emp}(\phi)=\frac{1}{N}\sum_{i=1}^{C}\sum_{j=1}^{n_{i}}\mathcal{L}(\phi(x^{i}_{j}),i)$ (1)
where $\mathcal{L}$ is the loss function. $n_{h}\gg n_{t}$ where $h<t$ Thus, misclassifying tail classes have much less risk compared with head classes. Generally, a considerable number of tail class samples are mislabeled as head classes. The straightforward solution involves increasing the importance or frequencies of tail classes. Re-weighting (RW) [4] or re-sampling (RS) [57, 1] can increase the risk of misclassification for tail classes by assigning larger weights or higher sampling rates to them. This, in turn, increases the risk of overfitting to tail classes. To expand the spatial distribution of tail classes and prevent overfitting, it is crucial to enhance the diversity of tail class samples. Unfortunately, obtaining more samples is generally infeasible. And Figure 1b shows that the tail classes are more prone to confusing with head classes. Typically, there are multiple instances of head and tail classes that are intermixed around the decision boundary. We leverage these intermixed instances by directly borrowing semantic information from head classes to augment tail classes. We have two purposes: i) increase the risk of misclassifying tail class samples to prevent the model from simply labeling them as head classes; ii) transfer the rich semantic information from head to tail classes, thereby improving the diversity of tail classes.
3.3 Fusing Head Features to Tail
Due to the imbalanced distribution of class samples, with the head classes being significantly more numerous than the tail classes, the latter tend to be crowded and underrepresented, leading to poor generalization performance on the test set of tail classes. As a large amount of tail class samples are misclassified as head classes, there may exist closet semantics between head and tail classes. To exploit this abundant closet semantic information, we fuse the features of head classes to the tail. This operation achieves the purpose of enriching the tail classes and expanding their feature space distribution. The fusion process is formulated as:
~$\tilde{\mathcal{F}}=\mathcal{M}_{p}\otimes\mathcal{F}^{t}+(\mathbf{I}-\mathcal{M}_{p})\otimes\mathcal{F}^{h},$ (2)
@@ 3.3 Fusing Head Features to Tail
where $\mathcal{M}_{p}$ is the mask combined with identity matrices $\mathbf{I}_{m}\in\mathcal{R}^{w_{F}\times h_{F}}$ and zero matrices $\mathbf{O}_{m}\in\mathcal{R}^{w_{F}\times h_{F}}$. The parameter $p$ represents the fusion ratio. Specially, the number of $\mathbf{I}_{m}$ and $\mathbf{O}_{m}$ is $d\cdot p$ and $d\cdot(1-p)$, respectively. $\tilde{\mathcal{F}}$ is then passed through a pooling layer and classifier to predict the corresponding logits $\mathbf{z}=[z_{1},z_{2},\cdots,z_{C}]$. Different loss functions such as CE loss, MisLAS [56], or GCL [27], to name a few, can be adopted. The backbone $\phi$ can choose the single model [13] as well as multi-expert model [47, 48]. We exploit the two-stage method and apply H2T in stage II.
Input: Training set, fusion ration $p$ ;
Output: Trained model;
1 Initialize the model $\phi$ randomly ;
2 for $iter=1$ to $E_{0}$ do
3       Sampling batches of data $(x,y)\sim\mathcal{T}^{I}$ from the instance-wise sampling data;
4       Obtain the feature map $\mathcal{F}=\phi_{\theta}(x)$ ;
5       Calculate logits $z=\mathbf{W}^{T}\mathbf{f}$ and loss $\mathcal{L}_{1}(x,y)$;
6       $\phi=\phi-\alpha\nabla_{\phi}\mathcal{L}_{1}((x,y);\phi)$.
7 end for
8for $iter=E_{0}+1$ to $E_{2}$ do
9       Sample batches of data $(x^{B},y^{B})\sim\mathcal{T}^{B}$ and $(x,y)\sim\mathcal{T}^{I}$;
10       Obtain feature maps $\mathcal{F}^{B}=\phi_{\theta}(x^{B})$ and $\mathcal{F}^{I}=\phi_{\theta}(x)$ ;
11       Fuse feature maps by Eq. 2 to obtain $\tilde{\mathcal{F}}$ ;
12       Input $\tilde{\mathcal{F}}$ to pooling layer to obtain $\tilde{\mathbf{f}}$, then calculate the logits by $\tilde{z}=\mathbf{W}^{T}\tilde{\mathbf{f}}$ and the loss by $\mathcal{L}_{2}(x^{B},y^{B})$ ;
13       Froze the parameters of representation learning $\phi^{r}$, and finetune the classifier parameters $\phi^{c}$: $\phi^{c}=\phi^{c}-\alpha\nabla_{\phi^{c}}\mathcal{L}_{2}((x^{B},y^{B});\phi^{c})$.
14 end for
@@ 10       Obtain feature maps $\mathcal{F}^{B}=\phi_{\theta}(x^{B})$ and $\mathcal{F}^{I}=\phi_{\theta}(x)$ ;
Visual recognition tasks often involve a vast number of categories. Manually screening fused categories can be a tedious and time-consuming process. We thereby design a simple strategy. We sample two versions of data: 1) class-balanced data $\mathcal{T}^{B}$ used to balance the ERM of each class, which is fed into the fused branch, and 2) instance-wise data $\mathcal{T}^{I}$ has a high probability to obtain head class samples, which is fed into the fusing branch. The sampling rates $p^{B}_{i}$ and $p^{I}_{i}$ for $\mathcal{T}^{B}$ and $\mathcal{T}^{I}$ are calculated by
$p^{B}_{i}=\frac{1}{C\cdot n_{i}},\>p^{I}_{i}=\frac{1}{N}.$ (3)
Balanced sampling data ensures that each class is sampled with equal probability $\frac{1}{C}$. The fewer the number of samples, the higher the probability of being resampled multiple times. The original data is distributed with a long tail, resulting in a head-biased dataset where the head classes have a higher probability of being sampled. Next, we can use the feature maps $\mathcal{F}^{B}$ and $\mathcal{F}^{I}$ obtained from $\mathcal{T}^{B}$ and $\mathcal{T}^{I}$ to substitute $\mathcal{F}^{t}$ and $\mathcal{F}^{h}$ in Eq. (2). By doing so, the features of the repeatedly sampled tail classes will be fused with the head class features with a higher probability. Different from the common practice of linearly combining a pair of inputs and their labels for augmentation, we only use the labels of the fused branch, namely the balance-sampled data, as ground truth to utilize the semantics of head classes to enrich tail classes. The proposed framework is illustrated in Figure 2 and Algorithm 1. The key steps are highlighted in rose-red arrow and text, respectively.

### v2

# Feature Fusion from Head to Tail for Long-Tailed Visual Recognition

@@ Abstract
The imbalanced distribution of long-tailed data presents a considerable challenge for deep learning models, as it causes them to prioritize the accurate classification of head classes but largely disregard tail classes. The biased decision boundary caused by inadequate semantic information in tail classes is one of the key factors contributing to their low recognition accuracy. To rectify this issue, we propose to augment tail classes by grafting the diverse semantic information from head classes, referred to as head-to-tail fusion (H2T). We replace a portion of feature maps from tail classes with those belonging to head classes. These fused features substantially enhance the diversity of tail classes. Both theoretical analysis and practical experimentation demonstrate that H2T can contribute to a more optimized solution for the decision boundary. We seamlessly integrate H2T in the classifier adjustment stage, making it a plug-and-play module. Its simplicity and ease of implementation allow for smooth integration with existing long-tailed recognition methods, facilitating a further performance boost. Extensive experiments on various long-tailed benchmarks demonstrate the effectiveness of the proposed H2T. The source code is available at https://github.com/Keke921/H2T.
@@ 1 Introduction
Deep models have shown remarkable capabilities in diverse visual recognition tasks (Hu et al. 2019; Cheung, Li, and Zou 2021; Minaee et al. 2022; Lan et al. 2024), yet their performance heavily relies on data that is evenly distributed across categories. In contrast, real-world data typically follows long-tailed distributions (Reed 2001; Liu et al. 2019; Zhang et al. 2023), which hinders model performance, especially on minority classes (referred to as tail classes in long-tailed data). This challenge has become one of the bottlenecks limiting the advancement of deep models. In recent times, a plethora of methods have emerged to address the problem of severe class imbalance in long-tailed data from different aspects, such as class-balancing methods (Chawla et al. 2002; Huang et al. 2016; Ren et al. 2018; Cui et al. 2019; Lin et al. 2020; Huang et al. 2020; Hong et al. 2021; Park et al. 2021), re-margining methods (Cao et al. 2019; Li, Cheung, and Lu 2022; Li, Cheung, and Hu 2022), data augmentation (Kim, Jeong, and Shin 2020; Wang et al. 2021a; Li et al. 2021; Park et al. 2022), and ensembling learning (Wang et al. 2021b; Li et al. 2022a; Jin et al. 2023). While these methods demonstrated the capability to yield robust predictions, they primarily focus on training a new model to acquire a relatively balanced embedding space and/or assembling multiple diverse networks. However, they overlook acquiring a more optimal classifier (determining the decision boundary), which is crucial for fully releasing the potential of the acquired backbone.
Figure 1: Comparison between decision boundaries produced by (a) existing methods and (b) proposed H2T.
This paper thereby aims to obtain a more effective classifier with the trained backbone. It is widely recognized that increasing class margins (Cao et al. 2019; Deng et al. 2019) and/or tightening the intra-class space (Wang et al. 2018, 2017) for the training set can improve the generalization performance of the model (Liu et al. 2021; Hu et al. 2023). However, while the clear margin is effective for balanced data, it may still have a biased decision boundary for long-tailed data. Take the binary case in the embedding space of the obtained backbone as an example, the decision boundary is usually the midline connecting the two class centroids. The head class has sufficiently sampled and its embedding space is fully occupied (Xiao et al. 2021). On the opposite, the tail class suffers from a scarcity of samples, leading to sparsely distributed semantic regions. The bias in the tail centroid persists due to head squeeze even if the existence of a clear margin. As a result, during the inference stage, a considerable number of samples that differ from the training set emerge, causing the erosion of well-defined margins. Consequently, numerous tail class samples are misclassified as head classes, as illustrated in Figure 1(a).
@@
To enrich the sparse tail class semantics and calibrate the bias in tail classes, we propose a direct and effective solution, named head-to-tail fusion (H2T), which grafts partial semantics from the head class on the tail class. In particular, the fusion operation in H2T involves the direct replacement of certain features of tail samples with those of head samples. This is based on the assumption that predictions on tail classes that have rare instances are easily affected by head classes that appear frequently. Transferring the head semantics can effectively fill the tail semantic area and the category overlap, which compels the decision boundary to shift closer to a more optimal one, as shown in Figure 1(b). Ultimately, the generalization performance of the tail classes can be improved. To streamline the fusion operation, we design an easy-to-implement strategy that takes full use of the obtained features without retraining the backbone. Specifically, we randomly substitute several channels in the feature maps of the class-balanced data with those of the head-biased data. This makes the feature maps of tail classes with a high probability of being fused with head classes. We apply H2T to the classifier tuning stage, and with just a few lines of code and a few epochs of training, remarkable performance improvements can be achieved. In addition, H2T does not alter the structure of the backbone or increase the network parameters. This characteristic makes it highly adaptable and compatible with various existing techniques, facilitating its seamless incorporation into diverse existing methods. The main contributions of this paper are summarized as follows:
@@
  • We propose a novel H2T that borrows information from head classes and applies it to augment tail classes without additional data or network parameters, which can release the potential of the well-trained backbone and obtain better decision boundaries.
@@
  • We devise a simple fusion strategy that partially substitutes the feature maps of an instance-wise sampling branch with those of a class-balanced sampling branch. This approach enables H2T to be implemented with ease and without requiring any modifications to the original backbone structure.
@@
  • Considerable performance improvement is achieved in both single and multi-expert models by integrating the H2T module into the existing structures. Extensive experiments on popular benchmarks validate the efficacy of the proposed method.
@@ 2 Related Work
This section mainly makes an overview of methods based on data augmentation which is the most relevant regime.
@@
Classical Augmentation: Classical augmentation methods including flip, rotate, crop, padding, and color jittering etc. (Szegedy et al. 2015; He et al. 2016) have been widely applied in deep models. On long-tailed data, these methods can improve model robustness and prevent overfitting to specific input patterns to a certain extent. Recently, AutoAugment (Cubuk et al. 2019), Fast AutoAugment (Lim et al. 2019), Population based augmentation (Ho et al. 2019) and Randaugment (Cubuk et al. 2020), etc. have been proposed to determine the best data augmentation strategy for a dataset automatically. These methods have been proven to be effective in increasing the classification accuracy for DNNs and also have demonstrated their efficacy on long-tailed data (Ren et al. 2020; Cui et al. 2021; Li et al. 2022b).
@@
MixUp Based Augmentation: MixUp (Zhang et al. 2018) is an effective data augmentation method that linearly combines a pair of row images and their labels. This method adds complexity control to the uncovered space in the data space through linearly interpolating discrete sample points, which reduces the model generalization error. Numerous studies (Zhang et al. 2021; Zhong et al. 2021; Li, Cheung, and Lu 2022) have experimentally proven that MixUp can significantly improve long-tailed recognition performance. CutMix (Yun et al. 2019), a variant of MixUp, augments data by cutting a patch from one input image and pasting it into another image within the training set. The ground truth labels are combined in proportion to the patch areas. Park et al. (Park et al. 2022) propose mixing the foreground patches from tail classes with the background images from head classes via CutMix, enabling the transfer of context-rich background information from head to tail.
@@
New Images Based Augmentation: In addition to the aforementioned methods, researchers have proposed alternative approaches that generate new informative samples for tail classes. One such method is M2m (Kim, Jeong, and Shin 2020), which translates samples from head classes to tail classes using a pretrained auxiliary classifier. Wang et al. (Wang et al. 2021a) utilize the learned encoded variation information from head classes to synthesize images for tail classes. Zada et al. (Zada, Benou, and Irani 2022) directly exploit pure noise images as tail class samples. Zhang et al. (Zhang et al. 2021) utilize class activation maps (CAM) (Zhou et al. 2016) to separate image foreground and background, followed by augmenting the foreground by flipping, rotating, and scaling, etc. The new samples are then generated by overlaying the augmented foreground on the unchanged background.
@@
Feature Level Augmentation: Data augmentation can also be performed in feature space. For example, Manifold MixUp (Verma et al. 2019) conducts the linear interpolation on the features of input images, which has been shown to yield better results in long-tail learning (Zhang et al. 2021; Zhou et al. 2020). Chu et al. (Chu et al. 2020) use CAM to decompose features into class-generic and class-specific components. Tail classes are then augmented by combining their class-specific components with the class-generic components. Another promising technique is to transfer knowledge from head classes to tail classes. For example, Yin et al. (Yin et al. 2019) and Liu et al. (Liu et al. 2020) enrich tail classes with the intra-class variance of head classes to balance the class distributions in feature space. Recently, meta-learning-based augmentation has been leveraged to address class imbalance. Liu et al. (Liu, Li, and Sun 2022) design a meta-embedding that uses a memory bank to enrich tail classes. MetaSAug (Li et al. 2021) learns class-wise covariance matrices by minimizing loss on a small balanced validation set and subsequently uses the learned matrices to semantically augment tail classes.
@@ 3.1 Preliminaries
We formally define the basic notations used in this paper before going into detail about our proposed method. We use $\{x,y\}$ to represent one input image and its corresponding label. The total number of classes is denoted by $C$, thus, we have $y\in\{0,1,\cdots,C-1\}$. The training set includes $N$ samples. Suppose that class $i$ has $n_{i}$ training samples. Then $N=\sum_{i}n_{i}$. For simplicity, we suppose $n_{0}\geq n_{1}\geq\cdots\geq n_{C-1}$. Feed $x$ into the backbone, we can obtain its feature maps before the last pooling layer, denoted as $\mathcal{F}=\left[F_{0},F_{1},\cdots,F_{d-1}\right]\in\mathbb{R}^{w_{F}\times h_{F}\times d}$, where $w_{F}$ and $h_{F}$ represent the width and height of the feature map, respectively. $d$ is the dimension of features in embedding space. The representation after the last pooling layer is ${f}\in\mathbb{R}^{d}$. The weight of the linear classifier is represented as $\textbf{W}=[w_{0},w_{1},\cdots,w_{C-1}]\in\mathbb{R}^{d\times C}$, where $w_{i}$ represents the classifier weight for class $i$. We use the subscripts $h,m$ and $t$ to indicate head, medium and tail classes, respectively. $z_{i}=w_{i}^{T}f$ represents the predicted logit of class $i$, where the subscript $i=y$ denotes the target logit and $i\neq y$ denotes the non-target logit.
@@ 3.2 Motivation
Head classes are significantly more numerous than tail classes in long-tailed data, resulting in a biased classifier, which leads to poor performance on the test set of tail classes. The straightforward solution involves increasing the importance or frequencies of tail classes  (Cao et al. 2019; Zhou et al. 2020; Ando and Huang 2017). They can increase the performance of tail classes, nevertheless, they also entail an elevated risk of overfitting. To adjust the decision boundary and prevent overfitting, it is crucial to enhance the diversity of tail class samples (Zhang et al. 2023; Yang et al. 2022). Unfortunately, directly obtaining more samples is generally infeasible. We can consider enriching the tail class by maximizing the utilization of existing features. Typically, the misclassified samples are unseen instances from the training set, and these instances tend to be interspersed in the vicinity of the decision boundary. This pattern provides us an opportunity to augment the diversity within tail classes and populate class margins with semantically meaningful samples. This augmentation process contributes to an optimal decision boundary. To achieve this, we simulate the potential unseen samples by directly borrowing semantic information from head classes to augment tail classes. By doing so, the rich semantic information can be transferred from well-represented head classes to tail classes.
3.3 Methodology: Fusing Head Features to Tail
We fuse the features of head classes to the tail to exploit the abundant closet semantic information. This operation can enrich the tail classes and expand their embedding space distribution. The fusion process is formulated as:
~$\tilde{\mathcal{F}}=\mathcal{M}_{p}\otimes\mathcal{F}_{t}+\overline{\mathcal{M}}_{p}\otimes\mathcal{F}_{h},$ (1)
@@ 3.3 Methodology: Fusing Head Features to Tail
where $\mathcal{M}_{p}$ is the mask stacked with multiple 1 matrices $\mathbf{I}_{M}=\mathbb{1}^{w_{F}\times h_{F}}$ and 0 matrices $\mathbf{O}_{M}=\mathbb{0}^{w_{F}\times h_{F}}$. The total number of all the 1 and 0 matrices is $d$. ¯$\overline{\mathcal{M}}{p}$ is the complement of $\mathcal{M}{p}$, that is, the indices of 0 matrices in ¯$\overline{\mathcal{M}}{p}$ correspond to the 1 matrix in $\mathcal{M}{p}$, and vice versa. The subscript $p$ of the mask matrices represents the fusion ratio. Specially, the number of $\mathbf{I}_{M}$ and $\mathbf{O}_{M}$ is $[d\times p]$ and $d-[d\times p]$, respectively. []delimited-[][$\cdot$][ ⋅ ] means rounding operation. $\tilde{\mathcal{F}}$ is then passed through a pooling layer and classifier to predict the corresponding logits $\mathbf{z}=[z_{0},z_{1},\cdots,z_{C-1}]$. Different loss functions, such as CE loss, MisLAS (Zhong et al. 2021), or GCL (Li, Cheung, and Lu 2022), to name a few, can be adopted. The backbone $\phi$ can be the single model (He et al. 2016) as well as the multi-expert model (Wang et al. 2021b; Xiang, Ding, and Han 2020). We exploit the two-stage training (Kang et al. 2020) and apply H2T in stage II.
@@
0:  Training set, fusion ration $p$
0:  Trained model
1:  Initialize the model $\phi$ randomly
2:  for $iter=1$ to $E_{0}$ do
3:     Sampling batches of data $(x,y)\sim\mathcal{T}^{I}$ from the instance-wise sampling data
4:     Obtain the feature map $\mathcal{F}=\phi_{\theta}(x)$
5:     Calculate logits $z=\mathbf{W}^{T}\mathbf{f}$ and loss $\mathcal{L}_{1}(x,y)$
6:     $\phi=\phi-\alpha\nabla_{\phi}\mathcal{L}_{1}((x,y);\phi)$.
7:  end for
8:  for $iter=E_{0}+1$ to $E_{2}$ do
9:     Sample batches of data $(x^{B},y^{B})\sim\mathcal{T}^{B}$ and $(x,y)\sim\mathcal{T}^{I}$
10:     Obtain feature maps $\mathcal{F}^{B}=\phi_{\theta}(x^{B})$ and $\mathcal{F}^{I}=\phi_{\theta}(x)$
11:     Fuse feature maps by Eq. 1 to obtain $\tilde{\mathcal{F}}$
12:     Input $\tilde{\mathcal{F}}$ to pooling layer to obtain $\tilde{\mathbf{f}}$, then calculate the logits by $\tilde{z}=\mathbf{W}^{T}\tilde{\mathbf{f}}$ and the loss by $\mathcal{L}_{2}(x^{B},y^{B})$
13:     Froze the parameters of representation learning $\phi^{r}$, and finetune the classifier parameters $\phi^{c}$: $\phi^{c}=\phi^{c}-\alpha\nabla_{\phi^{c}}\mathcal{L}_{2}((x^{B},y^{B});\phi^{c})$.
14:  end for
The selection of features to be fused poses a tedious task during training since visual recognition tasks often encompass a vast number of categories, and it cannot be guaranteed that each minibatch contains the required categories. We thereby devise a simple strategy to facilitate an easy-to-execute fusion process. This strategy involves sampling two versions of data: 1) class-balanced data $\mathcal{T}^{B}$ used to balance the empirical/structural risk minimization (ERM/SRM) of each class, which is fed into the fused branch, and 2) instance-wise data $\mathcal{T}^{I}$ has a high probability to obtain head class samples, which is fed into the fusing branch. The sampling rates $p^{B}_{i}$ and $p^{I}_{i}$ for class $i$ within the sets $\mathcal{T}^{B}$ and $\mathcal{T}^{I}$, respectively, are calculated by
$p^{B}_{i}=\frac{1}{C},\>p^{I}_{i}=\frac{n_{i}}{N}.$ (2)
Balanced sampling data ensures that each class is sampled with equal probability $\frac{1}{C}$. The fewer the number of samples, the higher the probability of being resampled multiple times. The instance-wise sampling branch deals with the original data distribution, resulting in head-biased data. Therefore, the head classes have a higher probability of being sampled. Next, we can use the feature maps $\mathcal{F}^{B}$ and $\mathcal{F}^{I}$ obtained from $\mathcal{T}^{B}$ and $\mathcal{T}^{I}$ to substitute $\mathcal{F}_{t}$ and $\mathcal{F}_{h}$ in Eq. (1). By doing so, the features of the repeatedly sampled tail classes will be fused with the head class features with a higher probability. In contrast to the common practice of linearly combining a pair of inputs and their labels for augmentation, we only use the labels of the fused branch, namely the balance-sampled data, as the ground truth. By adopting this method, the semantics of head classes are utilized to enrich tail classes. The proposed framework is illustrated in Figure 2 and Algorithm 1. The key steps are highlighted in rose-red.